OpenStudy (anonymous):

what is an inequality, why would I use it or try to solve it. I just found out that an exponet is not a number it is an abbreviation, so is an inequality something of the same kind of process

OpenStudy (anonymous):

An inequality gives bounds on some expression. An equality gives an exact solution. x = 4 means x is 4. \(x \le 4\) means x can be at most 4.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

So if someone tells you something like: Let w be the weight of a package. And the weight of a package is always between 2 and 5 pounds. You can translate that into an inequality: \[2\le w \le 5\]

OpenStudy (anonymous):

So far yes and you are the first person/now super hero to say such a good answer with awesome insight. Next, How would I know if a value is a solution for an inequality and inequality sign change when both sides are multiplied or divided by a negative. Why can't the textbook just tell us?

OpenStudy (anonymous):

Ok so a value is a solution to an inequality if it makes the inequality true.

OpenStudy (anonymous):

For example, which of the following are solutions to the inequality \(2 \le w \le 5\): 1, 2.5, 4, 7, -3

OpenStudy (anonymous):

the wieght example makes sense completely.

OpenStudy (anonymous):

2.4 and 4

OpenStudy (anonymous):

2.5, but yes.

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

You always change the direction of inequalities if you multiply or divide by a negative.

OpenStudy (anonymous):

Though I avoid this by adding/subtracting instead.

OpenStudy (anonymous):

Ok , so this is a rule for inequalities but what about in equations

OpenStudy (anonymous):

For example, lets say you had the inequality: \[1-w < 3\] You can subtract a 1 from both sides to get: \[-w < 2\] Then multiply both sides by -1 to get: \[w > -2\]

OpenStudy (anonymous):

Notice the direction changed.

OpenStudy (anonymous):

You don't have to worry about equalities because the equal sign is exact. It's not a range, so it doesn't have a direction.

OpenStudy (anonymous):

So we find that any w greater than -2 will be solutions to the original inequality (or any of the intermediate steps)

OpenStudy (anonymous):

For example, 3 is greater than -2, so if we plug in 3 for the original: \[1-3 \lt 3\] We get -2 < 3 which is true.

OpenStudy (anonymous):

But if we pick a w that's not greater than -2 (say -4 for example) \[1-(-4) < 3\] We get 5 < 3 which is false. So -4 does not satisfy this inequality.

OpenStudy (anonymous):

If you replace the equal sign of an equation with an inequality sign, is there ever a time when the same value will be a solution to both the equation and the inequality? yes, I see it changed is that because the bounds are now set to be anything higher than -1

OpenStudy (anonymous):

Only if the inequality is one of \[\le or \ge\] Because those inequalities allow for the possibility of it being equal.

OpenStudy (anonymous):

If it's strictly < or > then it cannot be =

OpenStudy (anonymous):

Any other questions?

OpenStudy (anonymous):

how did you know to write the expression -1(-4) ? 3 and why is is it a -5

OpenStudy (anonymous):

my computer will let me type a couple letters every few seconds

OpenStudy (anonymous):

9*6+3< 9*4 Yipee, this would be my first inequality

OpenStudy (anonymous):

Ok, well that one is just true or false.

OpenStudy (anonymous):

is 9*6 + 3 less than 9*4?

OpenStudy (anonymous):

Sometimes this site gets a bit laggy if the question-thread gets too long (like this one is). Perhaps we can continue this in a different question-thread.